# PAD-TRO: Projection-Augmented Diffusion for Direct Trajectory Optimization

### v2

@@ I INTRODUCTION
Recently, diffusion models have been widely adopted in computer vision tasks such as image and video generation [7, 15]. Due to its expressive power of modeling multi-modal distributions, several recent works have incorporated diffusion in trajectory optimization or control policy learning in robotics [3, 11, 20, 21, 19, 6, 18]. In particular, these diffusion-style methods can be divided into two categories: learning-based and model-based. Learning-based trajectory planning frameworks [3, 21, 6, 18] rely on expert demonstration data to train a noise prediction network. Then, trajectory generation is performed in a reverse diffusion process. They resemble the classic Denoising Diffusion Probabilistic Models (DDPM) [7]. In contrast, model-based diffusion frameworks [14, 11, 20] assume that we have some knowledge about the target distribution of optimal trajectories, and rely on Langevin sampling while estimating the score function, i.e., the gradient of the log-likelihood. For example, [14, 11, 20] assume that the target distribution of the optimal trajectories follows a Boltzmann distribution. Intuitively, the cost function is optimized by sampling from the corresponding Boltzmann distribution by annealing the temperature coefficient to zero [14]. Model-Based Diffusion (MBD) [14] solves the trajectory optimization problem by generating a sequence of control actions at each step of the reverse diffusion process. Then, the sequence of control actions in the last reverse diffusion step is applied to forward propagate the system dynamics, leading to a sequence of dynamically feasible states. In other words, MBD [14] is analogous to single shooting [8] in classic trajectory optimization. In contrast, Equality Constrained Diffusion for Direct Trajectory Optimization (DRAX) [11] generates both a sequence of states and actions via diffusion, and enforces soft dynamic feasibility by using an augmented Lagrangian term to penalize violation of dynamic feasibility. To the best of the authors’ knowledge, [11] is the only recent work that attempts to address nonlinear equality constraints in a model-based diffusion framework for direct trajectory optimization.
@@ II PROBLEM FORMULATION
where $p_{d}(X)=\prod_{t=1}^{T}\mathbf{1}(x_{t}=f(x_{t-1},u_{t-1}))$ denotes the dynamic feasibility condition, $p_{J}(X)\propto e^{-J(X)/\lambda}$ denotes optimality of the trajectory, and $p_{g}(X)\propto\prod_{t=1}^{T}\mathbf{1}\left(g_{t}(x_{t},u_{t})\leq 0\right)$ denotes inequality constraints satisfaction. The indicator function in $p_{d}(\cdot)$ and $p_{g}(\cdot)$ enforce the dynamic feasibility and obstacle avoidance of the trajectory. Indeed, the indicator function evaluates to 1 if the underlying constraints are satisfied; otherwise, it evaluates to 0. The cost of the trajectory $J$ is transformed into a Boltzmann distribution $p_{J}(\cdot)$ such that when the exponent is small, we get a value close to 1. The challenge of directly sampling from (2) arises from the high-dimensionality and sparsity of samples that strictly follow the target distribution $p_{0}(X)$. In particular, the dynamic constraints have zero measure. To tackle this challenge, our approach combines model-based diffusion with a gradient-free projection on an approximate reachable set. Before describing the algorithm in detail in Section IV, we discuss the relevant theoretical background of diffusion in the following section.
@@ IV-A Diffusion for direct sampling of state trajectory
To gradually generate samples that follow the target distribution $p_{0}(\cdot)$ (2), we propose the following approach built upon [14] by introducing a novel gradient-free projection scheme and a bi-level noise schedule. In [14], the noise schedule, i.e., the standard deviation of the Gaussian (12), $\sigma_{i}:=\sqrt{\frac{1-\bar{\alpha}_{i-1}}{\bar{\alpha}_{i-1}}}$ decays only along the diffusion horizon. Inspired by [20], we propose a bi-level noise schedule $\sigma_{i,t}$, where the noise not only varies along the diffusion horizon $i=N,N-1,\dots,1$, but also varies along the trajectory prediction horizon, $t=1,\dots,T$. In practice, we find that this allows for a better trade-off between exploration and reaching optimality in the reverse diffusion process. We define the novel noise schedule as follows:
$\sigma_{i,t}=\sqrt{\frac{1-\bar{\alpha}_{i-1}}{\bar{\alpha}_{i-1}}}\cdot\delta^{t},i\in[N,\dots,1],t\in[1,\dots,T].$ (10)
@@
$=\frac{\int\dfrac{-\tilde{x}^{i}-\sqrt{\bar{\alpha}_{i}}\,\tilde{x}^{0}}{1-\bar{\alpha}_{i}}\,p_{i|0}\!\left(\tilde{x}^{i}\mid\tilde{x}^{0}\right)p_{0}(\tilde{x}^{0})\,d\tilde{x}^{0}}{\int p_{i|0}\!\left(\tilde{x}^{i}\mid\tilde{x}^{0}\right)p_{0}(\tilde{x}^{0})\,d\tilde{x}^{0}}$ (19)
@@
$\nabla_{x^{i}}\log p_{i}(x^{i})=-\frac{\tilde{x}^{i}}{1-\bar{\alpha}_{i}}+\frac{\sqrt{\bar{\alpha}_{i}}}{1-\bar{\alpha}_{i}}\,\frac{\sum_{x\in\tilde{X}_{i}}p_{J}(x)p_{g}(x)}{\sum_{x\in\tilde{X}_{i}}p_{J}(x)p_{g}(x)},$ (23)
@@
Input : $\tilde{x}^{N}\in\mathbb{R}^{(T-1)\times n}\sim\mathcal{N}(0,I)$
@@ IV-B Dynamic feasibility via gradient-free projection
While we have introduced our direct trajectory sampling approach in Section IV-A that gradually generates samples from the target distribution $p_{0}(\cdot)\sim p_{J}(\cdot)p_{g}(\cdot)$, we have not yet accounted for dynamic feasibility $p_{d}(\cdot)$. Therefore, we augment our approach in Section IV-A by introducing a gradient-free projection to enforce dynamic feasibility recursively to complete our algorithm. Note that the predicted state sequence at reverse diffusion step $i$, i.e., $\tilde{x}^{i}_{t=1:T}$, is not yet dynamically feasible. Due to the recursive nature of dynamic feasibility, it is necessary to project the predicted state $\tilde{x}^{i}_{t+1}$ onto the reachable set of $\tilde{x}^{i}_{t}$. We denote such a reachable set as $\mathcal{H}(\tilde{x}^{i}_{t})$. We further denote $\mathcal{U}(\cdot)$ as a uniformly random distribution and denote the time-invariant admissible action set as $\mathcal{A}$. For a nonlinear system, we do not have access to the closed-form expression for such a reachable set $\mathcal{H}(\cdot)$.To address this challenge, we adopt a sampling approach: we uniformly draw a batch of $N_{p}$ action samples, with each sample consisting of all degrees of freedom of the action in $\mathbb{R}^{m}$. Thus, we obtain $\{U_{t}\}\in\mathbb{R}^{N_{p}\times m}\sim\mathcal{U}(\mathcal{A})$, where each action sample $u_{t}$ in $\{U_{t}\}$ has $m$ degrees of freedom. The admissible action set $\mathcal{A}:\{u^{\zeta}_{L}\leq u^{\zeta}\leq u^{\zeta}_{U}|\zeta=1,2,\dots,m\}$ refers to the allowable range of values between the lower bound $u_{L}$ and upper bound $u_{U}$ for each independent degree of freedom $\zeta$. Then, we forward propagate the predicted current state $\tilde{x}^{i}_{t}$ to obtain a batch of $N_{p}$ dynamically feasible samples $\{x^{i}_{t+1}\}=f(\tilde{x}^{i}_{t},u_{t})$, for all action samples $u_{t}\in\{U_{t}\}$. Next, we replace the predicted next state $\tilde{x}^{i}_{t+1}$ with the feasible sample in $\{x^{i}_{t+1}\}$ that is closest to $\tilde{x}^{i}_{t+1}$ in terms of the 2-norm distance. In summary, we project $\tilde{x}^{i}_{t+1}$ onto $\mathcal{H}(\tilde{x}^{i}_{t})$ as follows:
@@
and once we find the action $u^{*}_{t}$ that minimizes the distance $f(\tilde{x}^{i}_{t},u_{t})-\tilde{x}^{i}_{t+1}||^{2}_{2}$, our projected state (dynamically feasible) becomes $x^{i}_{t+1}=f(\tilde{x}^{i}_{t},u^{*}_{t})$. This process then repeats for all $t=0,1,\dots,T-2$. Our complete reverse diffusion process is formulated in Algorithm 1.

### v3

@@ I INTRODUCTION
Recently, diffusion models have been widely adopted in computer vision tasks such as image and video generation [7, 15]. Due to its expressive power of modeling multi-modal distributions, several recent works have incorporated diffusion in trajectory optimization or control policy learning in robotics [3, 11, 19, 20, 6, 18]. In particular, these diffusion-style methods can be divided into two categories: learning-based and model-based. Learning-based trajectory planning frameworks [3, 20, 6, 18] rely on expert demonstration data to train a noise prediction network. Then, trajectory generation is performed in a reverse diffusion process. They resemble the classic Denoising Diffusion Probabilistic Models (DDPM) [7]. In contrast, model-based diffusion frameworks [14, 11, 19] assume that we have some knowledge about the target distribution of optimal trajectories, and rely on Langevin sampling while estimating the score function, i.e., the gradient of the log-likelihood. For example, [14, 11, 19] assume that the target distribution of the optimal trajectories follows a Boltzmann distribution. Intuitively, the cost function is optimized by sampling from the corresponding Boltzmann distribution by annealing the temperature coefficient to zero [14]. Model-Based Diffusion (MBD) [14] solves the trajectory optimization problem by generating a sequence of control actions at each step of the reverse diffusion process. Then, the sequence of control actions in the last reverse diffusion step is applied to forward propagate the system dynamics, leading to a sequence of dynamically feasible states. In other words, MBD [14] is analogous to single shooting [8] in classic trajectory optimization. In contrast, Equality Constrained Diffusion for Direct Trajectory Optimization (DRAX) [11] generates both a sequence of states and actions via diffusion, and enforces soft dynamic feasibility by using an augmented Lagrangian term to penalize violation of dynamic feasibility. To the best of the authors’ knowledge, [11] is the only recent work that attempts to address nonlinear equality constraints in a model-based diffusion framework for direct trajectory optimization.
@@ II PROBLEM FORMULATION
where $p_{d}(X)\propto\prod_{t=1}^{T}\mathbf{1}(x_{t}=f(x_{t-1},u_{t-1}))$ denotes the dynamic feasibility condition, $p_{J}(X)\propto e^{-J(X)/\lambda}$ denotes optimality of the trajectory, and $p_{g}(X)\propto\prod_{t=1}^{T}\mathbf{1}\left(g_{t}(x_{t},u_{t})\leq 0\right)$ denotes inequality constraints satisfaction. The indicator function in $p_{d}(\cdot)$ and $p_{g}(\cdot)$ enforce the dynamic feasibility and obstacle avoidance of the trajectory. Indeed, the indicator function evaluates to 1 if the underlying constraints are satisfied; otherwise, it evaluates to 0. The cost of the trajectory $J$ is transformed into a Boltzmann distribution $p_{J}(\cdot)$ such that when the exponent is small, we get a value close to 1. The challenge of directly sampling from (2) arises from the high-dimensionality and sparsity of samples that strictly follow the target distribution $p_{0}(X)$. In particular, the dynamic constraints have zero measure. To tackle this challenge, our approach combines model-based diffusion with a gradient-free projection on an approximate reachable set. Before describing the algorithm in detail in Section IV, we discuss the relevant theoretical background of diffusion in the following section.
@@ IV-A Diffusion for direct sampling of state trajectory
To gradually generate samples that follow the target distribution $p_{0}(\cdot)$ (2), we propose the following approach built upon [14] by introducing a novel gradient-free projection scheme and a bi-level noise schedule. In [14], the noise schedule, i.e., the standard deviation of the Gaussian (12), $\sigma_{i}:=\sqrt{\frac{1-\bar{\alpha}_{i-1}}{\bar{\alpha}_{i-1}}}$ decays only along the diffusion horizon. Inspired by [19], we propose a bi-level noise schedule $\sigma_{i,t}$, where the noise not only varies along the diffusion horizon $i=N,N-1,\dots,1$, but also varies along the trajectory prediction horizon, $t=1,\dots,T$. In practice, we find that this allows for a better trade-off between exploration and reaching optimality in the reverse diffusion process.
We define the novel noise schedule as follows:
$\sigma_{i,t}=\sqrt{\frac{1-\bar{\alpha}_{i-1}}{\bar{\alpha}_{i-1}}}\cdot\delta^{t},i\in[N,\dots,1],t\in[1,\dots,T],$ (10)
@@
$=\frac{\int\dfrac{-\tilde{x}^{i}+\sqrt{\bar{\alpha}_{i}}\,\tilde{x}^{0}}{1-\bar{\alpha}_{i}}\,p_{i|0}\!\left(\tilde{x}^{i}\mid\tilde{x}^{0}\right)p_{0}(\tilde{x}^{0})\,d\tilde{x}^{0}}{\int p_{i|0}\!\left(\tilde{x}^{i}\mid\tilde{x}^{0}\right)p_{0}(\tilde{x}^{0})\,d\tilde{x}^{0}}$ (19)
@@
$\nabla_{x^{i}}\log p_{i}(x^{i})=-\frac{\tilde{x}^{i}}{1-\bar{\alpha}_{i}}+\frac{\sqrt{\bar{\alpha}_{i}}}{1-\bar{\alpha}_{i}}\,\frac{\sum_{x\in\tilde{X}_{i}}xp_{J}(x)p_{g}(x)}{\sum_{x\in\tilde{X}_{i}}p_{J}(x)p_{g}(x)},$ (23)
@@
Input : $\tilde{x}^{N}\in\mathbb{R}^{T\times n}\sim\mathcal{N}(0,I)$
@@ IV-B Dynamic feasibility via gradient-free projection
While we have introduced our direct trajectory sampling approach in Section IV-A that gradually generates samples from the target distribution $p_{0}(\cdot)\sim p_{J}(\cdot)p_{g}(\cdot)$, we have not yet accounted for dynamic feasibility $p_{d}(\cdot)$. Therefore, we augment our approach in Section IV-A by introducing a gradient-free projection to enforce dynamic feasibility recursively to complete our algorithm. Note that the predicted state sequence at reverse diffusion step $i$, i.e., $\tilde{x}^{i}_{t=1:T}$, is not yet dynamically feasible. Due to the recursive nature of dynamic feasibility, it is necessary to project the predicted state $\tilde{x}^{i}_{t+1}$ onto the reachable set of $\tilde{x}^{i}_{t}$. We denote such a reachable set as $\mathcal{H}(\tilde{x}^{i}_{t})$. We further denote $\mathcal{U}(\cdot)$ as a uniformly random distribution and denote the time-invariant admissible action set as $\mathcal{A}$. For a nonlinear system, we do not have access to the closed-form expression for such a reachable set $\mathcal{H}(\cdot)$.To address this challenge, we adopt a sampling approach: we uniformly draw a batch of $N_{p}$ action samples, with each sample consisting of all degrees of freedom of the action in $\mathbb{R}^{m}$. Thus, we obtain $\{U_{t}\}\in\mathbb{R}^{N_{p}\times m}\sim\mathcal{U}(\mathcal{A})$, where each action sample $u_{t}$ in $\{U_{t}\}$ has $m$ degrees of freedom. The admissible action set $\mathcal{A}:\{u^{\zeta}_{L}\leq u^{\zeta}\leq u^{\zeta}_{U}|\zeta=1,2,\dots,m\}$ refers to the allowable range of values between the lower bound $u_{L}$ and upper bound $u_{U}$ for each independent degree of freedom $\zeta$. Then, we forward propagate the predicted current state $\tilde{x}^{i}_{t}$ to obtain a batch of $N_{p}$ dynamically feasible samples by $\{x^{i}_{t+1}\}=f(\tilde{x}^{i}_{t},u_{t})$, for all action samples $u_{t}\in\{U_{t}\}$. Next, we replace the predicted next state $\tilde{x}^{i}_{t+1}$ with the feasible sample in $\{x^{i}_{t+1}\}$ that is closest to $\tilde{x}^{i}_{t+1}$ in terms of the 2-norm distance. In summary, we project $\tilde{x}^{i}_{t+1}$ onto $\mathcal{H}(\tilde{x}^{i}_{t})$ by first solving:
@@
Then, once we find the action $u^{*}_{t}$ that minimizes the distance $f(\tilde{x}^{i}_{t},u_{t})-\tilde{x}^{i}_{t+1}||^{2}_{2}$, our projected state (dynamically feasible) becomes $x^{i}_{t+1}=f(\tilde{x}^{i}_{t},u^{*}_{t})$, which completes the projection step. This process then repeats for all $t=0,1,\dots,T-2$. Our complete reverse diffusion process is formulated in Algorithm 1.